Feb. 14-16

– Three stores provided ~33.3 hours of luminosity
– Interlock personnel replace a bad MI card
– Controls personnel replace a bad Pbar card
– Tevatron cryo system quenches after store 8498 was terminated
– Booster personnel replaced power supply for low-level RF system
– Booster personnel repairs termination for Chopper
– Network Storm hits; Controls personnel located and isolated problem, but the storm affected many pieces of equipment
– Preaccelerator personnel switched from the H- Source to the I- Source
– Booster personnel worked on extraction issues
– DZero required an 8-hour access on Wednesday to repair detector

Read the Current Accelerator Update
Read the Early Bird Report
View the Tevatron Luminosity Charts